OpenStudy (anonymous):

solve using the elimination method. 3x+2y = 3 9x-8y=-2

OpenStudy (anonymous):

(3x+2y = 3) x -3 =( -9x-6y=-9) ( -9x-6y=-9) + (9x-8y=-2) -> (-14y = -11) y=11/14, x=10/21

OpenStudy (zepp):

Can you multiply the first equation by 4?

OpenStudy (anonymous):

U can multiply it by 3 to get the value for x

OpenStudy (zepp):

But the problem is, is you do that, you won't be able to cancel out the 9x because both 9x in both equations are positive.

OpenStudy (anonymous):

  Now substitute this value of y in any equation to get x